Benson Electronics currently produces the shipping containers it uses to deliver the electronics products it sells. The monthly cost of producing 9,000 containers follows:

Unit-level materials $ 5,400
Unit-level labor 6,300
Unit-level overhead 3,300
Product-level costs* 12,000
Allocated facility-level costs 26,900

*One-third of these costs can be avoided by purchasing the containers.

Russo Container Company has offered to sell comparable containers to Benson for $2.70 each.

Required

  1. Calculate the total relevant cost. Should Benson continue to make the containers?

  2. Benson could lease the space it currently uses in the manufacturing process. If leasing would produce $11,400 per month, calculate the total avoidable costs. Should Benson continue to make the containers?

a. Total relevant cost
Should Benson continue to make the containers?
b. Total avoidable cost
Should Benson continue to make the containers?
Answer #1

a. Relevant cost = 5400+6300+3300+12000**1/3 = $19000

Relevant cost per unit = 19000/9000 = $2.11 per unit

Benson should continue to make the container.

B. Total avoidable cost = 12000*1/3 + 11400 = $15400

Relevant cost = 5400+6300+3300+12000*1/3+11400 = $30400

Relevant cost per unit = 30400/9000 = $3.38 per unit

Benson should not continue to make the container.